Session: Sep 2000 (session 3)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Hartebeesthoek (Jonathan Quick at Sat Sep 23 02:50:41 METDST 2000) Success
Only LCP available, recorded in both channels just in case.

Onsala (A. Polatidis at Mon Sep 25 00:58:45 METDST 2000) Success
Recorded LINEAR polarization in both IF1 and IF2. The TSYS values taken before 14:32 in both IF1 and IF2 should be multiplied by 0.37 (apologies for this). After 14:32UT the values are OK.

Effelsberg (A. Kraus at Mon Sep 25 10:02:40 METDST 2000) Minor failures
LCP and RCP are exchanged! The receivers RCP-channel has much higher system temperature than LCP. More details on the 5cm-session page.

Torun (K.Borkowski at Mon Sep 25 10:31:46 METDST 2000) Success
Observations done according to schedule. It seems however that the frequency switchings were not effective when scheduled during passes (while tape moving). This could have been the case in other stations as well.

Jodrell_Bank (Alastair Gunn at Tue Sep 26 06:31:27 METDST 2000) Success
No known problems.

Noto (C.Trigilio at Tue Sep 26 14:12:16 METDST 2000) Failure
Failure of the tape recorder after first pass. Capstan motor out of order. Observation lost.